# Few-shot Object Localization

### v1

@@ IV-C Implementation Details
In our experiments, query images are uniformly resized to 512$\times$512, and the extracted query features are down-sampled to 128 $\times$ 128. At the same time, support features are uniformly cropped to 3$\times$3. The query features and support features are then uniformly projected to 256 dimensions. All experiments are implemented on a single NVIDIA RTX 3090. Since we mainly experiments on FSC-147 dataset, while support samples in this dataset are cropped from query images, the batch size is set to 1 to prevent negative impacts from other images. The number of support samples in all experiments is set to 1 to reflect the competitive performance of the FSOL model in the few-shot setting. In other word, all of our experiments are performed in one-shot setting. The model uses the Adam optimizer, with a learning rate of 0.00002, which drops by 0.25 every 80 epochs. All experiments are run for 200 epochs, and the epoch with the highest F1-score under $\sigma_{l}=10$ is taken as the optimal result.
IV-D Experiment Results
@@
Given that this paper introduces the first FSOL framework, there are no existing methods for comparison. To evaluate the FSOL model’s performance, we converted the current state-of-the-art few-shot counting method SafeCount into a localization model for comparison [16]. Firstly, we conducted experiments on the FSC-147 dataset, and the results showed that the performance of our model is superior to the most advanced few-shot counting method SafeCount (block = 1), while significantly saving computational resources and training time. To further verify the performance of the model, we also conducted positioning performance tests on three dense datasets. Compared with strong supervised models, our model still shows competitive performance under the few-shot setting. Additionally, we conducted ablation experiments to validate the effectiveness and necessity of the proposed DFA and SQ modules. These results indicate that our model has good generalization ability and potential, providing a strong baseline for the field of FSOL.
To explore the counting and positioning performance of the FSOL model, we chose to compare the performance with the most advanced few-shot object counting method SafeCount [16] on the FSC-147 dataset [37]. We firstly adapted SafeCount model into localization model and integrated it into our experimental pipeline. We then tested the models with 4 SafeCount blocks and 1 SafeCount block respectively. At the same time, we recorded the computational consumption of our proposed FSOL model, the baseline FSOL model which removed proposed DFA and SQ modules, and SafeCount models with different numbers of blocks. We recorded the parameter quantity, GFLOPs, and the average time it takes for processing single image for 4 models respectively. For fairness, the average results of ten experiments are taken for each set of experiment. Specifically, when calculating the processing time of single image, we first take the average time required for the model to process 100 images, then conduct ten experiments and take the average to get the final result. Table 1 records all the above experimental results.
TABLE I: Quantitative results on FSC-147 dataset.
@@
Params(M) GFLOPs(G) Speed(ms) MAE MSE F1($\sigma=5$) F1($\sigma=10$)
SafeCount(1 block) 53.22 474.86 184.3 24.4 70.84 52.76 67.94
@@
FSOL Baseline 48.44 113.59 27.00 23.05 77.11 43.41 62.57
FSOL + DFA + SQ 48.57 116.95 75.03 20.33 68.28 53.40 70.00
@@
From the results, it’s evident that the FSOL model maintains competitive localization and counting performance compared to the SafeCount model while significantly reducing the required parameter quantity, computational resources, and computation time. Specifically, compared to the 1 block and 4 blocks SafeCount models, the FSOL model reduces the parameter quantity by about 5M and 20M, respectively. The GFLOPs are about one-fifth and one-seventh, respectively, and the computation time is about half and one-fourth, respectively. Under these conditions, when $\sigma_{l}=10$, the F1-score increases by 2% and decreases by 5%, respectively. In such a trade-off, the FSOL model, compared to the 1 block SafeCount model, not only improves both counting and localization performance but also saves computational resources, indicating the proposed FSOL model’s excellent performance and great potential.
@@
To investigate the model’s performance in dense object prediction tasks, we also conducted experiments on dense datasets ShangHaiTech PartA, ShangHaiTech PartB, and CARPK [21, 41]. In addition, we recorded the performance of the baseline model to obtain the performance improvement results brought by the two modules we proposed. At the same time, we selected five strong supervised localization models FIDTM [18], CLTR [42], GMS [43], IIM [44], and STEERER [45] framework for comparison with the FSOL model and FSOL Baseline model to demonstrate the significant localization performance improvement brought by our model. The parameters and experimental settings of the mentioned models can be found in their original papers. Among them, FIDTM and CLTR use the low threshold $\sigma_{s}$ of 4 and high threshold $\sigma_{l}$ of 8. In this experiment, the FSOL model uses thresholds consistent with these. GMS, IIM, and STEERER did not provide low threshold results, the high threshold they use is $\sigma_{l}=\sqrt{w^{2}+h^{2}}/2$, where $w$ and $h$ represent the width and height of the sample. For IIM, for consistency with FSOL, we choose point annotation model results during comparison. The results in Table 2 and Table 3 show that the proposed FSOL model shows competitive performance compared to strong supervised models even in few-shot settings.
TABLE II: Quantitative results on ShangHaiTech A dataset.
Supervised Models Threshold=$\sigma_{s}$ Threshold=$\sigma_{l}$
@@
TABLE III: Quantitative results on ShangHaiTech B dataset.
Supervised Models Threshold=$\sigma_{s}$ Threshold=$\sigma_{l}$
@@
We also conducted model tests on the remote sensing dataset CARPK, using experimental settings of $\sigma_{s}=5,\sigma_{l}=10$. From the experimental results in Table 4, it can be seen that when $\sigma_{l}=10$, the F1-score increased from 84.76% to 93.46%. It can be concluded that the DFA and SQ module greatly improved the overall performance.
TABLE IV: Quantitative results on CARPK dataset.
Models Threshold=$\sigma_{s}$ Threshold=$\sigma_{l}$
@@ IV-E Ablation Study
To verify the effectiveness of the proposed modules, we conducted extensive ablation studies on the FSC-147 dataset. We conducted experiments by removing Self-query (SQ) module, Dual-path Feature Augmentation (DFA) module, Deformable Convolution (DC) used in DFA, and Cross Central Difference Convolution (CCD-C) used in DFA, respectively. Table 5 records the quantitative analysis of the SQ experiment results, while Table 6 presents the quantitative analysis of the remaining experiments. The results show that the modules we proposed have a positive impact on localization performance.
IV-E1 Self-query
To demonstrate the effectiveness of the Self-query (SQ) module, we removed the SQ module and compared the experimental results with FSOL model. As mentioned earlier, the similarity map $S\in\mathbb{R}^{1\times 256\times 128\times 128}$ input to the SQ module and the output similarity map $S_{SQ}\in\mathbb{R}^{1\times 256\times 128\times 128}$ share the same dimensions and size. Therefore, to explore the performance improvement brought by the SQ module, we simply removed it. Table 5 shows that removing SQ significantly decreases the localization performance on both the validation set and the test set. Specifically, when the low threshold $\sigma_{s}=5$ is selected, which means the judgment of positive and negative samples is stricter, on the validation set and test set, the F1-score decreased by 2.86% and 2.07% respectively, AP decreased by 4.51% and 2.51% respectively, and AR decreased by 1.35% and 1.55% respectively; when the high threshold $\sigma_{l}=10$ is used, on the validation set and test set, the F1-score decreased by 2.79% and 2.66% respectively, AP decreased by 4.94% and 3.31% respectively, and AR decreased by 0.83% and 1.89% respectively. This shows that the proposed SQ module can significantly contribute to FSOL.
TABLE V: Quantitative results between FSOL model and SQ-removed FSOL model.
Self_Query Validation Set Test Set
@@ IV-E1 Self-query
TABLE VI: Quantitative results between FSOL model, DFA-removed FSOL model and CCD-C or DC replaced models.
@@
IV-E2 Dual-path Feature Augmentation
To demonstrate the effectiveness of the Dual-path Feature Augmentation (DFA) module, we uniformly replace Deformable Convolution (DC) and Cross Central Difference Convolution (CCD-C) with vanilla 2D convolution. This ensures that the query feature $F_{Q}\in\mathbb{R}^{1\times 256\times 128\times 128}$ and support feature $F_{S}\in\mathbb{R}^{1\times 256\times 3\times 3}$ are consistent in spatial dimensions with the output features of the DFA module. Table 6 shows that when the DFA module is removed or modified with vanilla convolution, the localization performance significantly decreases on both the validation set and the test set. We focus on the performance change when F1-score on the validation set and test set when the high threshold $\sigma_{l}=10$ is selected. Specifically, when the DFA module is removed, that is, the CCD-C and DC used for feature enhancement are all replaced with vanilla convolution, the F1-score decreases by 5.25% and 4.76% respectively; when the CCD-C in DFA is replaced with vanilla convolution, the F1-score decreases by 0.74% and 0.83% respectively; as for DC is replaced, the F1-score decreases by 1.74% and 3.02% respectively. This shows that the DFA module not only stimulates the potential of CCD-C convolution and DC in the two branches respectively, but also positively contributes to the feature fusion of the two branches.
@@ V Future Work
The FSOL framework we proposed provides a high-performance baseline for the FSOL task, serving as a reference and platform for future work in this area. However, while the Dual-path Feature Augmentation (DFA) and Self-query (SQ) module we proposed have brought competitive performance improvements, their structure is still very simple and needs further improvement. In future work, we plan to make improvements from the following aspects:
V-A SQ Module Upgrade
Along the idea of self-query, we will further design adaptive module structures, so that the FSOL framework can align with different styles of query images, further exert the self-query ability of the proposed SQ model, and thus enable the FSOL framework to have cross-domain matching capabilities.
V-B DFA Module Upgrade
The current DFA module simply uses two branches to enhance deformation information and gradient information respectively, and then uses 3D convolution to fuse these two pieces of information. Future work could focus on exploring other feature information that helps sample matching, and then fuse them in a more interpretable and more elegant way.
V-C Model Architecture Upgrade
Future research could explore more complex architectures, such as Transformer or the latest Mamba architecture [46]. These architectures may better capture complex patterns and dependencies in the data, which may further improve the accuracy of object localization.
V-D Task Upgrade
Future work could focus on diversifying the modalities of query prompts, such as using natural language instead of support prompt images to search for corresponding information in query images, thereby achieving the leap from FSOL to Zero-shot Object Localization, and making contributions in a larger field.
@@ VI Conclusion
The FSOL framework we proposed not only defines a new domain in few-shot learning but also provides a high-performance, robust few-shot benchmark for it. The Dual-path Feature Augmentation (DFA) module we proposed plays a key role in understanding, comparing, and fusing features. By guiding the model, it enables the generation of more accurate sample-query image similarity maps, thereby helping the model understand the differences in samples at the feature level and enhancing the model’s ability to search for given samples in query images. The Self-query (SQ) module we proposed introduces the query image’s information into the similarity map’s optimization using the concept of self-query, making the sample-query image similarity map more realistic. As a result, our FSOL model can be more targeted to understand the differences within the class samples, thereby identifying more intra-class samples, and thus increase the accuracy and generalization of the model. The experimental results on sparse and dense datasets reveal the powerful performance and great application value of proposed FSOL model.

### v3

@@ IV-C Implementation Details
In our experiments, query images are uniformly resized to $512\times 512$, and the extracted query features are down-sampled to $128\times 128$. At the same time, support features are uniformly cropped to $3\times 3$. The query features and support features are then uniformly projected to 256 dimensions. All experiments are implemented on a single NVIDIA RTX 3090. Since we mainly experiments on FSC-147 dataset, while support samples in this dataset are cropped from query images, the batch size is set to 1 to prevent negative impacts from other images. The model uses the Adam optimizer, with a learning rate of 0.00002, which drops by 0.25 every 80 epochs. All experiments are run for 200 epochs, and the epoch with the highest F1-score under $\sigma_{l}=10$ is taken as the optimal result.
IV-D Experiment Results and Analysis
@@
Since this paper introduces the first FSOL framework, there are no existing methods available for comparison. To evaluate the FSOL model’s performance, we adapt the current state-of-the-art few-shot counting method, SafeCount [16], into a localization model for comparison. Initially, experiments are conducted on the FSC-147 dataset, revealing superior performance of our model compared to the most advanced few-shot counting method, SafeCount (block = 1), while also significantly reducing computational resources and training time. Furthermore, localization performance tests are carried out on three dense datasets to further validate the model’s performance. Even when compared with robust supervised models, our model demonstrates competitive performance under the few-shot setting. Additionally, ablation experiments are conducted to verify the effectiveness and necessity of the DFA and SQ modules. These findings suggest that our model exhibits strong generalization ability and potential, establishing a solid baseline for the field of FSOL.
To assess the counting and localization performance of the FSOL model, we compare it with SafeCount on the FSC-147 dataset. Initially, we integrate SafeCount into our pipeline as a localization model, testing it with both 4 and 1 SafeCount blocks under 1-shot setting. We then record the computational consumption of our FSOL model, the baseline model (without DFA and SQ modules), and SafeCount models with varying block counts, including parameters, GFLOPs, and average processing time per image. For fairness, we average results over ten experiments, especially for processing time per image, where we calculate the average time required for the model to process 100 images across ten experiments. Furthermore, we demonstrate the performances of our FSOL model in both 2-shot and 3-shot settings to showcase its generalization capabilities, while also highlighting its ability to prevent overfitting in the 1-shot setting. Table I summarizes these findings.
TABLE I: Quantitative comparison results on FSC-147 dataset. The comparison metrics encompass the model’s computational cost, as well as its localization and counting performance.
@@
[…]
SafeCount(1 block) 53.22 474.86 184.30 24.4 70.84 52.76 67.94
@@
Baseline 48.44 113.59 27.00 23.05 77.11 43.41 62.57
1-shot FSOL 48.57 117.38 60.40 20.33 68.28 53.40 70.00
2-shot FSOL 48.57 118.60 85.27 20.78 72.29 54.15 70.59
3-shot FSOL 48.57 119.82 105.74 19.85 68.25 54.75 71.36
@@
The results demonstrate that the FSOL model maintains competitive localization and counting performance compared to the SafeCount model, while significantly reducing the required parameter quantity, computational resources, and computation speed. SafeCount utilizes transformer-based multi-head self-attention and direct query feature addition to achieve weighted feature aggregation and residual feature fusion. Nonetheless, within our FSOL pipeline, the weighted feature aggregation is substituted with 3D convolution of information-enhanced features, while residual fusion is achieved through selective Self-Query weighting guided by the query featu re. Consequently, our FSOL pipeline not only achieves superior localization performance but also reduces the model’s complexity. Specifically, compared to the 1-block and 4-block SafeCount models, the FSOL model reduces the parameter quantity by approximately 5M and 20M, respectively. The GFLOPs are approximately one-fifth and one-seventh, respectively, and the computation speed is approximately half and one-fourth, respectively. Under these conditions, when $\sigma_{l}=10$, the F1-score increases by 2% and decreases by 5%, respectively. In such a trade-off, the FSOL model not only improves both counting and localization performance compared to the 1-block SafeCount model but also saves computational resources, highlighting the excellent performance and potential of the FSOL model.
Furthermore, the counting and localization performances of our FSOL model in 1-shot, 2-shot, and 3-shot settings underscore its robustness in few-shot scenarios. In other words, our FSOL model is capable of effectively learning from the information within the support samples and generalizing to novel samples, even when the number of provided support samples is limited. Moreover, our FSOL model achieves its optimal performance in the 3-shot setting. This suggests that our FSOL model can exploit an increased number of given samples to enhance its localization capabilities, while also demonstrating superior adaptability, thereby achieving competitive performance across various experimental settings. Further, we visualize the localization results of our FSOL model on the FSC-147 dataset, as depicted in Fig. 6. Through qualitative analysis, it is evident that our FSOL model is capable of handling both dense and sparse objects with competitive performance. Additionally, the multi-shot results underscore the learning ability and robustness of our FSOL model.
Figure 6: Qualitative localization results on the FSC-147 dataset are visualized. The MAE and F1-score represent the mean values calculated from three images under various experimental settings. The FSOL model achieves a lower MAE and a higher F1-score when provided with an increased number of support samples.
@@
To investigate the model’s performance in dense object localization tasks, we conduct experiments on dense datasets ShangHaiTech PartA, ShangHaiTech PartB, and CARPK [21, 39]. Additionally, we record the performance of the baseline model to assess the performance improvement brought by the two modules. We select five strong supervised localization models: FIDTM [18], CLTR [40], GMS [41], IIM [42], and the STEERER [43] for comparison with the FSOL model and Baseline model. This helps demonstrate the significant localization performance improvement brought by our model. FIDTM and CLTR use a low threshold $\sigma_{s}$ of 4 and a high threshold $\sigma_{l}$ of 8. In this experiment, the FSOL model uses thresholds consistent with these. GMS, IIM, and STEERER do not provide low threshold results; they use a high threshold of $\sigma_{l}=\sqrt{w^{2}+h^{2}}/2$, where $w$ and $h$ represent the width and height of the sample. For consistency with FSOL, we choose point annotation model results for IIM during comparison. The results in Table II and Table III show that the FSOL model demonstrates competitive performance compared to strong supervised models, even in few-shot settings.
TABLE II: Quantitative comparison results on ShangHaiTech A dataset.
Supervised Methods Threshold=$\sigma_{s}$ Threshold=$\sigma_{l}$
@@
TABLE III: Quantitative comparison results on ShangHaiTech B dataset.
Supervised Methods Threshold=$\sigma_{s}$ Threshold=$\sigma_{l}$
@@
We also conduct model tests on the remote sensing dataset CARPK, using experimental settings of $\sigma_{s}=5$ and $\sigma_{l}=10$. From the experimental results in Table IV, it can be seen that when $\sigma_{l}=10$, the F1-score increases from 84.76% to 93.46%. This indicates that the DFA and SQ module greatly improve the overall performance.
TABLE IV: Quantitative comparison results on CARPK dataset.
Methods Threshold=$\sigma_{s}$ Threshold=$\sigma_{l}$
@@ IV-E Ablation Study
To verify the effectiveness of our modules, we conduct extensive ablation studies on the FSC-147 dataset. We experiment by removing the Self Query (SQ) module, Dual-path Feature Augmentation (DFA) module, Deformable Convolution (DC) used within DFA, and Cross Central Difference Convolution (CCD-C) used in DFA, respectively. Table V records the quantitative analysis of the SQ experiment results, while Table VI presents the quantitative analysis of the remaining experiments. The results indicate that these modules have a positive impact on localization performance.
IV-E1 Self Query Ablation
To illustrate the effectiveness of the SQ module, we remove it and compare the experimental results with the FSOL model. Table V demonstrates that removing SQ notably decreases the localization performance on both the validation set and the test set. Specifically, when the low threshold $\sigma_{s}=5$ is selected, indicating stricter judgment of positive and negative samples, the F1-score decreases by 2.86% and 2.07% respectively on the validation and test sets. Similarly, the AP decreases by 4.51% and 2.51% respectively, and AR decreases by 1.35% and 1.55% respectively. When the high threshold $\sigma_{l}=10$ is used, the F1-score decreases by 2.79% and 2.66% respectively on the validation and test sets. The AP decreases by 4.94% and 3.31% respectively, and AR decreases by 0.83% and 1.89% respectively. This underscores the significant contribution of the SQ module to FSOL.
TABLE V: Quantitative comparison results between FSOL and SQ-removed models.
Self query Validation Set Test Set
@@ IV-E1 Self Query Ablation
TABLE VI: Quantitative comparison results among FSOL, DFA-removed and CCD-C or DC replaced models.
@@
IV-E2 Dual-path Feature Augmentation Ablation
To demonstrate the effectiveness of the DFA module, we uniformly replace Deformable Convolution (DC) and Cross Central Difference Convolution (CCD-C) with vanilla 2D convolution. This ensures that the query feature $F_{Q}\in\mathbb{R}^{1\times 256\times 128\times 128}$ and support feature $F_{S}\in\mathbb{R}^{1\times 256\times 3\times 3}$ are consistent in spatial dimensions with the output features of the DFA module. Table VI shows that when the DFA module is removed or modified with vanilla convolution, the localization performance significantly decreases on both the validation set and the test set. We focus on the performance change when F1-score on the validation set and test set when the high threshold $\sigma_{l}=10$ is selected. Specifically, when the DFA module is removed, that is, the CCD-C and DC used for feature enhancement are all replaced with vanilla convolution, the F1-score decreases by 5.25% and 4.76% respectively; when the CCD-C in DFA is replaced with vanilla convolution, the F1-score decreases by 0.74% and 0.83% respectively; as for DC is replaced, the F1-score decreases by 1.74% and 3.02% respectively. This shows that the DFA module not only stimulates the potential of CCD-C convolution and DC in the two branches respectively, but also positively contributes to the feature fusion of the two branches.
Figure 7: Visualization of the localization results under different support samples. The figure showcases various selections of support samples and visualizes the corresponding localization results. It is clear that the choice of support samples greatly affects the model’s localization performance.
@@ V Future Work
Our FSOL framework establishes a strong baseline for the FSOL task, providing a solid foundation for future research in this field. While our Dual-path Feature Augmentation (DFA) and Self Query (SQ) modules have demonstrated competitive performance enhancements, their simplicity suggests potential for further refinement. In future work, we aim to improve our approach in the following ways:
V-1 Support Samples Selection
In our 1-shot FSOL pipeline, the selection of the support sample from the available dataset significantly impacts localization performance. Fig. 7 illustrates varying object localization outcomes based on different selected support samples. Looking ahead, our goal is to develop a more adaptive and robust localization method capable of extracting deeper semantic information from the provided support samples, thereby ensuring precise and reliable localization.
V-2 Model Architecture Upgrade
Future research could delve into exploring more advanced architectures, such as Transformer or the latest Mamba architecture [44]. These architectures hold promise in capturing intricate patterns and dependencies within the data, potentially leading to improved accuracy in object localization.
V-3 Task Upgrade
In future research, there could be an emphasis on diversifying the modalities of query prompts. For example, exploring the use of natural language instead of support prompt images to search for corresponding information in query images. This approach could potentially facilitate the transition from few-shot object localization to zero-shot object localization, thereby making significant contributions to a broader field.
@@ VI Conclusion
This paper defines the FSOL task for the first time and proposes a high-performance baseline framework. The framework consists of two main modules: The DFA module ingeniously strengthens the association between support samples and query images through its dual-branch design, mitigating significant intra-class appearance gaps and addressing object omission challenges; The SQ module integrates the distribution information of the query image into the similarity map, resulting in a substantial enhancement of the model’s performance. Experimental results on both sparse and dense datasets demonstrate the robust performance and significant application potential of our FSOL model. With this high-performance FSOL framework, the model can effectively handle unseen target categories given only a few labeled samples, greatly facilitating future studies in FSOL task as well as expanding the scope of object localization applications.